OpenStudy (anonymous):

An acute angle θ is in a right triangle with sin θ =2/3 . What is the value of cot θ?

OpenStudy (ghazi):

 now by using pythagoras theorem find the third side

OpenStudy (anonymous):

9+4=13

OpenStudy (ghazi):

let's say third side is x then x=\[\sqrt{3^2-2^2}\]

OpenStudy (ghazi):

x=\[\sqrt{5}\]

OpenStudy (anonymous):

and then

OpenStudy (ghazi):

then cot theta= base/perpendicular= \[\sqrt{5}/2\]

OpenStudy (anonymous):

lifesaver !!

OpenStudy (ghazi):

i guess !! by the way according to pythagoras theorem 3^2= 2^2+sqrt5^2